# Controlling Large Electric Vehicle Charging Stations via User Behavior Modeling and Stochastic Programming

### v3

@@ IV-C Uncertainty distribution approximation
  • Kwh requests. The regressor $r_{1}$ is used to obtain the kWh request associated with a session start (transition from “inactive” to “active” state). It approximates $\mathbf{k}$, the function that returns the kWh requests given $(g,l,\tilde{i})$, as defined in IV-B2. Note that $\delta^{i}_{t}$ , the announced parking time associated with a session start on slot $i$ at time $t$ is not predicted. This approach yielded better results than the one based on session end prediction. Consequently, the remaining number of time steps before the announced disconnection of EV connected to slot $i$, $m^{i}_{t}$ is set to $\infty$- ∞ as a flag.
@@
  • Session ends. The second classifier is used to obtain the transition probabilities from the “active” to “inactive” state, and thus, the session end uncertainty samples $(q_{t_{0}},\tilde{q}_{t_{0}+1},...,\tilde{q}_{t_{0}+R})$. It approximates: $p^{i}_{1\to 0}=\mathbb{P}(\mathbf{O}(t)=0|\mathbf{O}(t-1)=1,g=g^{i}_{t-1},l=l_{t-1},\tilde{i}=i)$
Algorithm 1 is used to generate the uncertainty samples using $c_{1},c_{2}$ and $r_{1}$; its working principle is illustrated in Fig. 1.
@@
This modeling shows numerous improvements over the one introduced in [12]. Regarding the end-of-session times, our novel model allows for correcting the tendency of users to disconnect their vehicles before the announced end of parking time, while [12] was assuming this announced time to be the effective end-of-session time. In [12], the controller could plan on charging the EVs on the very last time steps of the announced parking period, which would lead to customer dissatisfaction if the users disconnected their EVs earlier. By allowing our model to predict session ends to happen before the expected time, this risk is reduced.
Moreover, as opposed to the heuristic proposed in [12] consisting in setting the load of free slots to constant values to simulate future charging sessions; planning the sessions start, sessions end, and kWh request allows for more accurate modeling of the future EVCS load.
@@
Figure 1: Working principle of a single iteration of the scenario generation algorithm on a single charging slot. The text in red indicates the corresponding lines in the code of Algorithm 1
Algorithm 1 Uncertainty sample drawing: Draw a sample of the future uncertainties $({w}_{t_{0}},\tilde{w}_{t_{0}+1},...,\tilde{w}_{t_{0}+R})$ from $\mathbb{P}(\mathbf{W}_{t_{0}},\mathbf{W}_{t_{0}+1},\ldots,\mathbf{W}_{t_{0}+R}|\mathbf{H}_{t_{0}}=h_{t_{0}})$
1:function GET_SAMPLE($x_{t_{0}},h_{t_{0}}$)
2:    for $t\in[t_{0}+1...t_{0}+R]$ do
3:       $g_{t-1},m_{t-1},l_{t-1},a_{t-1},\delta_{t-1}\leftarrow\text{GET\_FEATURES($h_{t-1})$}$triangleright Compute features from uncertainty history
4:       for $i,~{}i<n$ do
5:          $o^{i}_{t}\leftarrow o^{i}_{t-1}$triangleright set current state to be the previous one by default
6:          if $o^{i}_{t}=1$ then $\triangleright$ if slot $i$ was in active state…
7:             if $m^{i}_{t-1}=0$ then
8:                $o^{i}_{t}\leftarrow 0$triangleright end session if announced parking time elapsed
9:             elsenormal-$\triangleright$ if annouced parking time has not elapsed…
10:                $p\leftarrow c_{2}(g_{t-1},l_{t-1},i)$
11:                if RANDOM.RAND() < $p$ then $\triangleright$ end session with probability p
12:                   $o^{i}_{t}\leftarrow 0$
13:                end if
14:             end if
15:             $h_{t}\leftarrow\text{UPDATE\_HISTORY($g_{t-1},m_{t-1},l_{t-1},a_{t-1},\delta_{t-1},o_{t}$)}$triangleright update history
16:          end if
17:       end for
18:
19:       $g_{t},m_{t},l_{t},a_{t},\delta_{t}\leftarrow\text{GET\_FEATURES($h_{t})$}$triangleright Update features from uncertainty history
20:       for $i,~{}i<n$ do
21:          if $o^{i}_{t}=0$ then $\triangleright$ if slot was in inactive state (or has just transitioned to inactive state)…
22:             $p\leftarrow c_{1}(g_{t},l_{t},i)$
23:             if RANDOM.RAND() < $p$ then $\triangleright$ start session with probability p
24:                $o^{i}_{t}\leftarrow 1$
25:                $k^{i}_{t}\leftarrow r_{1}(g_{t},l_{t},i)$triangleright get kWh request
26:                $m^{i}_{t}\leftarrow-\infty$triangleright set announced parking time duration to $\infty$- ∞ (as a flag)
27:             end if
28:          end if
29:       end for
30:
31:       $h_{t}\leftarrow\text{UPDATE\_HISTORY($g_{t},m_{t},l_{t},a_{t},\delta_{t},o_{t}$)}$
32:    end for
33:    $({w}_{t_{0}},\tilde{w}_{t_{0}+1},...,\tilde{w}_{t_{0}+R})\leftarrow h_{t_{o}:t_{o}+R}$
34:return $({w}_{t_{0}},\tilde{w}_{t_{0}+1},...,\tilde{w}_{t_{0}+R})$
35:end function
@@ IV-D Uncertainty forecasts for MPC
To obtain different uncertainty samples (used with 2S), the switches between states are done with probability obtained using $c_{1}$ and $c_{2}$, allowing different scenarios to be obtained for the same state conditions. As introduced in IV-A2, our MPC algorithm uses an uncertainty forecast instead of relying on a collection of uncertainty samples. In this context, the switch between states is done in a deterministic way, i.e. only if the probability of state switch (obtained using $c_{1}$ and $c_{2}$) is greater than the probability of staying in the same state. In practice, this means that lines 13, and 25 are modified to compare p with 0.5 instead of a random value.
@@ IV-E1 User request-based MPC (R-MPC)
Assuming that slot $i$ is “active” and that the last request $r_{t_{0}}^{i}=(k_{t_{0}}^{i},\delta_{t_{0}}^{i})$, we get:
[…]
This baseline is the most straightforward MPC approach that can be deployed using the available user data. However, it does not anticipate the end of charging sessions by disconnections happening before the announced parking time elapses and does not account for potential future charging sessions that could increase the load in future time steps, as it only uses an average load. This becomes problematic if the EVCS delays the charge of the current EVs on those time steps — e.g., because of off-peak prices, for example — as it results in unsatisfied charging requests.
@@ V-A1 Datasets
We use the Caltech Adaptive charging network dataset [17]. This dataset contains charging session logs with duration, connection and disconnection time, kWh request, announced parking duration, and slot identifiers.
The portion used for training covers 83 days and contains 1123 sessions across $n=32$ slots. Each slot has at least 20 training sessions. The one used for testing covers 22 days, and contains 352 sessions over the same number of slots.
The dataset is pre-processed to obtain fixed interval data points of $\Delta t=15$ min. Connection and disconnection times are rounded to the nearest quarter-hour, and the charging session information is encoded using the uncertainty variables introduced in section II.
Unsatisfiable user requests are rounded down to the largest amount of energy that can be provided in the span of the announced parking time. This means that all requests can be satisfied if the users let their EV connected to the slot for the parking duration that they anticipated. However, this does not always happen in practice. These early disconnections are responsible for P-MPC not being able to reach 100% satisfaction.
@@ V-A2 EVCS settings
The EVCS parameters are of an existing EVCS station. We set the maximum slot power to 12kW. Each slot can thus deliver an energy amount equal to $\bar{e}=3$kWh in the span of one time step (15 minutes). We further set the subscribed power limit to 8% of the total nominal capacity of the EVCS, i.e. $\bar{c}=0.08\cdot n\cdot\bar{e}=7.68$kWh per time step. This corresponds to two times the average EVCS load. The electricity prices are set to 0.102€  during off-peak hours (i.e. from 00:00 to 06:00, 09:00 to 11:00, 13:00 to 17:00, and 21:00 to 00:00) and to 0.153€  during peak hours. Similarly, the threshold overrun penalty is set to $\xi=14.31$€ for every time step spent above the threshold, as per the “Tarif Jaune” rate from France’s main electric utility company EDF [18]. Finally, the battery charge efficiency of the connected EVs is set to $\eta=0.91$
V-B Implementation specifics
@@
V-C Evaluation metrics
The algorithms are compared based on the following criteria:
  • Electricity costs: Total cost of electricity, including the contract threshold overrun penalties.
@@ V-C Evaluation metrics
    • Filling rate: What percentage of the user request is provided on average by the end of the session.
    • Full satisfaction rate: Percentage of requests that are fully satisfied (an energy amount corresponding to the initial request has been provided) at the end of session time.
@@
$\alpha$ Alg. Objective Electricity Filling Full satisfaction Electricity cost relative Filling rate relative Full satisfaction relative
@@
500 2S 99056 12991 69% 60% 16% 3% -1%
500 MPC 98600 12392 68% 59% 10% 1% -2%
500 R-MPC 177245 32942 70% 52% 194% 4% -14%
1000 2S 139141 17286 77% 66% 3% -4% -9%
1000 MPC 139401 17156 77% 65% -4% -5% -10%
1000 R-MPC 222683 34519 72% 53% 93% -10% -26%
5000 2S 404811 43094 85% 67% 3% -6% -16%
5000 MPC 445859 38484 81% 61% -8% -10% -23%
5000 R-MPC 564019 37859 73% 55% -9% -19% -31%
50000 2S 1744329 138893 92% 83% 122% -1% -2%
50000 MPC 2097777 117179 89% 78% 87% -4% -7%
50000 R-MPC 4214607 45470 74% 54% -27% -21% -36%
V-D Results
We report in Table I the results obtained for different $\alpha$, the weight given to the satisfaction component of the objective function:
@@ V-D Results
As the value of $\alpha$ increases $\blacksquare$, the two satisfaction metrics both increase for all algorithms, and our 2S and MPC algorithms show very similar results to P-MPC. For $\alpha=5000$, 2S is only 3% more costly but achieves 94% of the performance of P-MPC in terms of filling rate, and 84% in terms of full satisfaction rate. These values are respectively 16% and 23% higher than the ones obtained with our industry-grade baseline R-MPC.
As the value of $\alpha$ gets very large $\blacksquare$, the results obtained by 2S are close to the ones obtained with P-MPC in terms of satisfaction, while those obtained with MPC are slightly lower. With $\alpha=10^{5}$, The filling rates of 2S and P-MPC are almost identical (92% for 2S against 93% for P-MPC), and the full satisfaction rates are equal (84% for both algorithms). 2S improves the filling rate obtained using R-MPC, our industry-grade baseline by 24%, and the full satisfaction rate by 55%. The electricity costs are significantly higher for both MPC and 2S (respectively 101% and 136% higher than with P-MPC), as customer unsatisfaction is much more penalized than electricity costs. Note that the fact that P-MPC does not reach 100% of satisfaction rate is because of early disconnections of sessions that could could otherwise been satisfied.
Fig. 3 demonstrates that 2S is more robust than MPC in terms of customer satisfaction, as it achieves higher values for both satisfaction metrics. This can be attributed to the fact that the control action is derived from a greater number of uncertainty scenarios, increasing the likelihood of anticipating early session ends caused by early disconnections. Consequently, the controller is incentivized to initiate charging earlier and at a higher charging rate in these situations. Additionally, 2S achieves higher satisfaction scores than MPC for the same electricity costs, positioning itself as the best-performing algorithm that can be realistically implemented. Moreover, 2S significantly improves the results obtained by MPC for every value of $\alpha$. The improvements of 2S over R-MPC are particularly significant for large values of $\alpha$, for which the full satisfaction rate of 2S is 55% higher than the one obtained with R-MPC.
Finally, we show in Fig. 2 the distribution of the percentage of the initial request that was left to be provided at disconnection time, for the different algorithms with $\alpha=5000$. R-MPC shows the largest values, as it uses the announced end of session time without correcting the user bias, so that many EVs get disconnected before they receive the requested energy amount. In average for $\alpha=5000$, P-MPC2S, MPC and R-MPC respectively took 164ms, 1563ms, 556ms, and 193ms per time step on a modern laptop with an Intel i7 1185G7.
@@ VI Conclusion
In this paper, we presented an Electric Vehicle Charging Station (EVCS) model that considers numerous real-world constraints, including slot power limitations, penalties for contract threshold overruns, and early disconnections of EVs. We proposed a mathematical formulation for the optimal control problem under uncertainty and implement two Multi-Stage Stochastic Programming approaches: Model Predictive Control and Two-Stage Stochastic Programming. These approaches leverage the information provided by EV users. The problem complexity arises from uncertainties in charging session start and end times, as well as the energy requested by users. This necessitates predicting future charging session times and their energy demand. Early disconnections pose an additional challenge, as customers may disconnect their EVs before the initially announced time, further complicating charge planning. To address this, we introduced a user behavior model based on a sojourn time-dependent stochastic process that considers the duration of the charging session, allowing for cost reduction while maintaining customer satisfaction.
Our algorithms were compared to two baselines: a Model Predictive Control algorithm with a perfect forecast of the future uncertainties (P-MPC), representing a theoretical optimal controller, and an industry-grade Model Predictive Control algorithm mentioned in [12]. Simulations were conducted using a real-world dataset spanning 22 days. The results demonstrated the advantages of our methods, which incorporate user behavior modeling and leverage user-provided information. Particularly, the Two-Stage Stochastic Programming method proves robust against early disconnections by considering a greater number of uncertainty scenarios for optimization. Our Two-Stage algorithm, prioritizing user satisfaction over electricity cost, performs comparably to the optimal baseline, outperforming our previous algorithm from [12] by 24% and 55% in the two user satisfaction metrics, respectively. The algorithm that strikes the best balance between cost and user satisfaction shows a 3% relative cost difference (slightly higher) compared to the optimal baseline, while achieving 94% and 84% of the user satisfaction performance of the optimal baseline in the two metrics used.
In the near future, we plan on studying the performance of our algorithm on charging sessions for which no end-of-parking time was specified by the user. Another major addition to this work would be to add a second optimization layer that would take advantage of the idle time (the period when an EV is still connected, although the announced end of parking time has elapsed) to complete the charge of the EVs that could not be charged on time. Finally, future work will also consist in incorporating EVCS control into the control of a larger microgrid that additionally features battery storage systems and PV production. This transition aims to move beyond solving local optimization problems towards establishing broader control systems and will require combining EV behavior modeling, battery charge scheduling [22], and netload modeling [15].
Acknowledgments
The authors sincerely thank Claude Lepape, Jean-Christophe Alais, and Alejandro Yousef Da Silva for their valuable input and support.

### v4

@@ IV-C2 Kwh requests
The regressor $r_{1}$ is used to obtain the kWh request associated with a session start (transition from “inactive” to “active” state). It approximates $\mathbf{k}$, the function that returns the kWh requests given $(g,l,\tilde{i})$.
IV-C3 Session ends
@@
The second classifier is used to obtain the transition probabilities from the “active” to “inactive” state, and thus, the session end uncertainty samples $(q_{t_{0}},\tilde{q}_{t_{0}+1},...,\tilde{q}_{t_{0}+R})$. It approximates: $p^{i}_{1\to 0}=\mathbb{P}(\mathbf{O}(t)=0|\mathbf{O}(t-1)=1,g=g^{i}_{t-1},l=l_{t-1},\tilde{i}=i)$
@@
The working principle of the algorithm used to generate the uncertainty samples using $c_{1},c_{2}$ and $r_{1}$ is illustrated in Fig. 1 . Regarding the end-of-session times, our novel model allows for correcting the tendency of users to disconnect their vehicles before the announced end of parking time, while [10] was assuming this announced time to be the effective end-of-session time. In [10], the controller could plan on charging the EVs on the very last time steps of the announced parking period, leading to customer dissatisfaction if the users disconnected their EVs earlier. By allowing our model to predict session ends to happen before the expected time, this risk is reduced. Moreover, as opposed to the heuristic proposed in [10] consisting in setting the load of free slots to constant values to simulate future charging sessions; planning the sessions start, sessions end, and kWh request allows for more accurate modeling of the future EVCS load.
@@
Figure 1: Working principle of a single iteration of the scenario generation algorithm on a single charging slot.
@@ IV-D Uncertainty forecasts for MPC
To obtain different uncertainty samples (used with 2S), the switches between states are done with probability obtained using $c_{1}$ and $c_{2}$, allowing different scenarios to be obtained for the same state conditions.
@@ IV-E1 User request-based MPC (R-MPC)
This baseline is the most straightforward MPC approach that can be deployed using the available user data. However, it does not anticipate the end of charging sessions by disconnections before the announced parking time elapses and does not account for potential future charging sessions that could increase the load in future time steps. This becomes problematic if the EVCS delays the charge of the current EVs on those time steps — e.g., because of off-peak prices — as it results in unsatisfied charging requests.
@@ V-A1 Datasets
We use the Caltech Adaptive charging network dataset [15]. The training covers 83 days and contains 1123 sessions across $n=32$ slots. Each slot has at least 20 training sessions. The test set covers 22 days and 352 sessions over the same number of slots. The dataset is pre-processed to obtain fixed interval data points of $\Delta t=15$ min. Unsatisfiable user requests are rounded down to the largest amount of energy that can be provided in the span of the announced parking time. This means that all requests can be satisfied if the users let their EV connected to the slot for the parking duration that they anticipated. However, this does not always happen in practice. These early disconnections are responsible for P-MPC not being able to reach 100% satisfaction.
@@ V-A2 EVCS settings
The EVCS parameters are of an existing EVCS station. We set the maximum slot power to 12kW. Each slot can thus deliver an energy amount equal to $\bar{e}=3$kWh in the span of one time step (15 minutes). We further set the subscribed power limit to 8% of the total nominal capacity of the EVCS, i.e. $\bar{c}=0.08\cdot n\cdot\bar{e}=7.68$kWh per time step. This corresponds to two times the average EVCS load. The electricity prices are set to 0.102€  during off-peak hours (i.e. from 00:00 to 06:00, 09:00 to 11:00, 13:00 to 17:00, and 21:00 to 00:00) and to 0.153€ during peak hours. Similarly, the threshold overrun penalty is set to $\xi=14.31$€ for every time step spent above the threshold, as per the “Tarif Jaune” rate from France’s main electric utility company EDF [16]. Finally, the battery charge efficiency of the connected EVs is $\eta=0.91$
V-A3 Implementation specifics
@@
V-A4 Evaluation Metrics
The algorithms are compared based on cost and two customer satisfaction criteria:
  • Electricity Costs: Total cost, including contract threshold penalties.
@@ V-A4 Evaluation Metrics
    • Filling Rate: Average percentage of user request provided by session end.
    • Full Satisfaction Rate: Percentage of requests fully satisfied by session end.
@@
{tblr}
cells = c, cell11 = r=2, cell12 = r=2, cell13 = r=2, cell35 = PeriwinkleGray, cell36 = PeriwinkleGray, cell37 = PeriwinkleGray, cell38 = PeriwinkleGray, cell39 = PeriwinkleGray, cell45 = PeriwinkleGray, cell46 = PeriwinkleGray, cell47 = PeriwinkleGray, cell48 = PeriwinkleGray, cell49 = PeriwinkleGray, cell65 = PastelGreen, cell66 = PastelGreen, cell67 = PastelGreen, cell68 = PastelGreen, cell69 = PastelGreen, cell75 = PastelGreen, cell76 = PastelGreen, cell77 = PastelGreen, cell78 = PastelGreen, cell79 = PastelGreen, cell95 = PastelGreen, cell96 = PastelGreen, cell97 = PastelGreen, cell98 = PastelGreen, cell99 = PastelGreen, cell105 = PastelGreen, cell106 = PastelGreen, cell107 = PastelGreen, cell108 = PastelGreen, cell109 = PastelGreen, cell125 = PetiteOrchid, cell126 = PetiteOrchid, cell127 = PetiteOrchid, cell128 = PetiteOrchid, cell129 = PetiteOrchid, cell135 = PetiteOrchid, cell136 = PetiteOrchid, cell137 = PetiteOrchid, cell138 = PetiteOrchid, cell139 = PetiteOrchid, $\alpha$ Alg. Electricity Filling Full satisfaction Electricity cost relative Filling rate relative Full satisfaction relative
@@
500 2S 12991 69% 60% 16% 3% -1%
500 MPC 12392 68% 59% 10% 1% -2%
500 R-MPC 32942 70% 52% 194% 4% -14%
1000 2S 17286 77% 66% 3% -4% -9%
1000 MPC 17156 77% 65% -4% -5% -10%
1000 R-MPC 34519 72% 53% 93% -10% -26%
5000 2S 43094 85% 67% 3% -6% -16%
5000 MPC 38484 81% 61% -8% -10% -23%
5000 R-MPC 37859 73% 55% -9% -19% -31%
50000 2S 138893 92% 83% 122% -1% -2%
50000 MPC 117179 89% 78% 87% -4% -7%
50000 R-MPC 45470 74% 54% -27% -21% -36%
V-B Results
We present the results in Table I for varying $\alpha$, representing the weight assigned to customer satisfaction:
@@ V-B Results
As the value of $\alpha$ increases $\blacksquare$, the satisfaction metrics both increase for all algorithms, and our 2S and MPC algorithms show similar results to P-MPC. For $\alpha=5000$, 2S is only 3% more costly but achieves 94% of the performance of P-MPC in terms of filling rate, and 84% in terms of full satisfaction rate. These values are respectively 16% and 23% higher than the ones obtained with our industry-grade baseline R-MPC.
As the value of $\alpha$ gets very large $\blacksquare$, the results obtained by 2S are close to the ones obtained with P-MPC in terms of satisfaction, while those obtained with MPC are slightly lower. With $\alpha=10^{5}$, The filling rates of 2S and P-MPC are almost identical (92% for 2S against 93% for P-MPC), and the full satisfaction rates are equal. 2S improves the filling rate obtained using R-MPC, our industry-grade baseline by 24%, and the full satisfaction rate by 55%. The electricity costs are significantly higher for both MPC and 2S (respectively 101% and 136% higher than with P-MPC), as customer unsatisfaction is much more penalized than electricity costs. Note that the fact that P-MPC does not reach 100% of satisfaction rate is because of early disconnections of sessions that could have otherwise been satisfied.
Fig. 3 illustrates 2S’s superior robustness in customer satisfaction compared to MPC, attributed to a broader range of uncertainty scenarios. This enables better anticipation of early session ends, resulting in earlier and faster charging actions. 2S consistently outperforms MPC across all satisfaction metrics for equivalent electricity costs, positioning it as the most viable implementation choice. Additionally, 2S significantly outperforms MPC, particularly for high $\alpha$ values.
Finally, Fig. 2 displays the distribution of unfulfilled energy requests at disconnection time for $\alpha=5000$. R-MPC exhibits the highest values, reflecting its reliance on user-provided session end times without correction. On average, for $\alpha=5000$, P-MPC, 2S, MPC, and R-MPC each took 164ms, 1563ms, 556ms, and 193ms per time step on an Intel i7 1185G7 laptop.
@@ VI Conclusion
In this paper, we presented an Electric Vehicle Charging Station model that considers real-world constraints, including slot power limitations, penalties for contract threshold overruns, and early disconnections of EVs. We proposed a mathematical formulation for the optimal control problem under uncertainty and implemented two Multi-Stage Stochastic Programming approaches: Model Predictive Control (MPC), and Two-Stage Stochastic Programming (2S). The problem complexity arises from uncertainties in charging session start and end times, as well as the energy requested by users. This necessitates predicting future charging session times and their energy demand. Early disconnections pose an additional challenge, as customers may disconnect their EVs before the initially announced time, further complicating charge planning. To address this, we introduced a user behavior model based on a sojourn time-dependent stochastic process that considers the duration of the charging session, allowing for cost reduction while maintaining customer satisfaction.
Our algorithms were compared to two baselines: a Model Predictive Control algorithm with a perfect forecast of the future uncertainties (P-MPC), representing a theoretical optimal controller, and an industry-grade Model Predictive Control algorithm (R-MPC), mentioned in [10]. Simulations were conducted using a real-world dataset spanning 22 days. The results demonstrated the advantages of our methods, which incorporate user behavior modeling and leverage user-provided information. Particularly, 2S proves robust against early disconnections by considering a greater number of uncertainty scenarios for optimization. The version of 2S that prioritizes user satisfaction over electricity cost performs comparably to the optimal baseline P-MPC, outperforming our previous algorithm from [10], R-MPC, by 24% and 55% in the two user satisfaction metrics, respectively. The version of 2S that strikes the best balance between cost and user satisfaction shows a 3% relative cost difference (slightly higher) compared to the optimal baseline P-MPC, while achieving 94% and 84% of the user satisfaction performance of P-MPC in the two metrics used.